If the reflective Index of a medium is 1.5cm and the angle of Incidence=60 calculate the angle of refraction

1
Expert's answer
2022-03-14T09:45:52-0400

"n=\frac{\sin a}{\sin b},"

"b=\arcsin\frac{\sin a}n=35°."